OpenStudy (anonymous):

15/12.9 Without calculator how do you do this???

hero (hero):

@eloiseq, \[\frac{15}{12.9} = \frac{15}{12 + \frac{9}{10}}\] right?

OpenStudy (anonymous):

agreed

hero (hero):

Remember 12.9 is read as "12 and nine-tenths" which is the same as \(12 + \dfrac{9}{10}\)

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):